Can you please tell me if this is right? I'm not sure if there is any other way to simplify/complete this problem, or if this is it??? Thanks sooo much!
Find the coordinates of the point whose images under the translation (x,y)-->(x-3, y+4) are the following...
(h, k)
So would the right answer simply be: (h-3, k+4)????
Thank you so much in advance!!!

Answer by stanbon(75887)

Can you please tell me if this is right? I'm not sure if there is any other way to simplify/complete this problem, or if this is it??? Thanks sooo much!
Find the coordinates of the point whose images under the translation
(x,y)-->(x-3, y+4) are the following...
(h, k)
So would the right answer simply be: (h-3, k+4)????
-------------------------------------------------
x Coordinate:
The translation subtracts 3 from the input value
so you want x-3 = h
Solving for x you get x = h+3
------------------------------------
y Coordinate:
The traslation adds 4 to the input value
so you want k = y+4
Solving for y you get y = k-4
----------------------------------
Checking:
(h+3,k-4)---->(h+3-3,k-4+4) = (h,k)
=======================================
Cheers,
Stan H.